OpenStudy (anonymous):

Find the center and radius of x^2+y^2+2x-10y+10=0

OpenStudy (anonymous):

"complete the square" \[x^2+2x+y^2-10y=-10\] \[(x+1)^2+(y-5)^2=-10+1+25=16\] now read off the answer from the general form \[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (anonymous):

so the center would be 1 and -5 and the radius would be 16

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[r^2=16,r=4\]

OpenStudy (anonymous):

\[(x-h)^2+(y-k)^2=(x+1)^2+(y-5)^2\] \[h=-1,k=5\]

OpenStudy (anonymous):

ok thanks for explaining.

OpenStudy (anonymous):

yw